# DeepSeek vs. ChatGPT: A Comparative Study for Scientific Computing and Scientific Machine Learning Tasks

### v1

@@ 2.2 Scientific Machine Learning
In this section, we test the LLMs on a range of tasks in scientific machine learning, in both image recognition and physics-informed machine learning methods.
@@ 2.2.1 MNIST digits prediction
Many problems in science and engineering require analyzing image data and classification. The MNIST [46] dataset is chosen as a test problem because it is a well-established benchmark for evaluating image recognition models. All DeepSeek and ChatGPT models are asked with the following prompt:
@@
Table 5: Comparison of models generated by DeekSeek and ChatGPT for learning the MNIST dataset. “Convolution($n$)” means a 2D convolution layer with $n$ channels, “Dense($n$)” means a fully connected layer with $n$ neurons, and “Dropout($p$)” means dropout with probability $p$.
DeekSeek V3 DeepSeek R1 ChatGPT 4o ChatGPT 03-mini-high
Architecture Convolution(32) Convolution(32) Convolution(32) Convolution(32)
MaxPooling MaxPooling MaxPooling Convolution(64)
Convolution(64) Convolution(64) Convolution(64) MaxPooling
MaxPooling MaxPooling MaxPooling Dense(128)
Convolution(64) Dense(128) Dropout(0.5) Dense(10)
Dense(64) Dropout(0.5) Dense(128)
Dense(10) Dense(10) Dense(10)
Eoochs 5 20 10 10
Time (s) 21.30 21.23 21.85 21.18
Testing accuracy 98.96% 99.11% 99.31% 98.84%
Early stopping N/A Yes N/A N/A
Note that the DeepSeek R1 model implements an early stopping criterion, which may stop the training of the model early according to changes in the validation loss to prevent overfitting, which was activated in the 10-th epoch. Additionally, both DeepSeek R1 and ChatGPT 4o implement dropout with a probability of 0.5. All models achieved comparable training time and high testing accuracy.
@@ 2.2.2 Physics-informed neural networks
In this section, we will test their performance by solving the above equation with physics-informed neural networks (PINNs)[38], a framework used to solve PDEs with the power of neural networks. The Poisson equation we want to solve here is
$\begin{split}-\Delta u(x,y)=1,(x,y)\in[-1,1]^{2}/[0,1]^{2}\end{split}$
with zero Dirichlet boundary conditions. Therefore, the question that we ask the four models is:
@@
Table 6: Network architectures, hyperparameters, and $L_{2}$ prediction errors for four PINN models solving the Poisson equation. The architecture [2,20,20,20,1] represents a fully connected neural network with four hidden layers, each containing 20 neurons, an input dimension of 2, and an output dimension of 1.
DeepSeek V3 ChatGPT 4o DeepSeek R1 ChatGPT o3-mini-high
Network [2, 20, 20, 20, 1] [2, 64, 64, 64, 1] [2, 50, 50, 50, 50, 1] [2, 64, 64, 64,1]
Training data 1000 5000 512 200 (every epoch)
Epoch 5000 5000 10000 5000
Training time (s) 7.90 15.23 13.80 1740.00
Activation function Tanh Tanh Tanh Tanh
$L_{2}$ error 929% 160% 54.7% 5.3%
@@
Figure 5: Predicted solutions from different models for the Poisson equation.
We observe that only ChatGPT o3-mini-high gives the right result, which outperforms other models. Moreover, compared to non-reasoning models, DeepSeek-R1 can realize the boundary of the problem domain is L-shaped and thus generate corresponding boundary points. The non-reasoning model can not realize this and still train the model in the original problem domain. These behaviors demonstrate the reasoning model can have better generalization abilities and better responses to user-specified requests. The detailed parameter settings are listed in Table 6. It is clear that the reasoning model can achieve much better accuracy compared to the non-reasoning model. Furthermore, as for the training time, ChatGPT o3-mini-high takes over 1000 seconds to train the model as it needs to generate samples every training epoch, but it achieves the best performance.
@@ 2.2.3 DeepONet for learning the antiderivative operator
    DeepSeek V3: Input functions: simply generated point-wise as uniformly random numbers in numpy. The functions are then numerically integrated using cumulative sums (cumsum) in numpy.
@@
    DeepSeek R1: Data are randomly generated from the following 4 types:
@@
      Input function: A 6th-degree polynomial with uniformly random coefficients ranging from $[-2,2]$. Antiderivative: calculated analytically.
@@
      Input function: $u(x)=A\sin(\omega x)+B\cos(\omega x)$, where $A,B$ are chosen randomly from $[-2,2]$, and $\omega$ from $[1,5]$. Antiderivative: calculated analytically.
@@
      Input function: $u(x)=A\exp(Bx)$, where $A,B$ are chosen randomly from $[-1,1]$. Antiderivative: calculated analytically.
@@
      Input function: a mixture of 1, 2, and 3 given as $u(x)=c_{1}x^{2}+c_{2}\sin(\omega x)+c_{3}\exp(x)$, where $c_{i}$ are chosen randomly from $[-1,1]$ and $\omega$ from $[1,5]$. Antiderivative: calculated analytically.
@@
    ChatGPT 4o: Input functions: 5-th degree polynomial without constant term: $u(x)=c_{1}x+\cdots c_{5}x^{5}$, where the coefficients $c_{1}$ are chosen randomly from $[-1,1]$. Antiderivative: calculated analytically.
@@
    ChatGPT o3-mini-high: Input functions: Fourier series with 3 modes $u(x)=\sum_{i=1}^{3}A_{i}\sin(2\pi ix)+B_{i}\cos(2\pi ix)$ where $A_{i},B_{i}$ are chosen randomly from $[-1,1]$. Antiderivative: calculated numerically using cumulative trapezoidal rule with scipy.integrate.cumtrapz.
The model architectures are summarized in Table 7. Some bugs were fixed from the code generated by the LLMs:
@@
    ChatGPT 4o’s code incorrectly sets the trunk net input dimension to 100 (the number of evaluation points) instead of 1 (the dimension of the dependent variable). This mistake was also manually corrected.
Table 7: Network architectures, hyperparameters, and $L_{2}$ prediction errors for the four DeepONet models learning the antiderivative operator. The network architecture notation follows the same convention as in Table 6.
DeekSeek V3 DeepSeek R1 ChatGPT 4o ChatGPT o3-mini-high
Branch Net [100, 128, 128, 128] [100,100,100,50] [64, 64, 64, 64, 64] [100,100,100,100]
Trunk Net [100, 128, 128, 128] [100,100,100,50] [64, 64, 64, 64, 64] [100,100,100,100]
Num of sensors 100 100 100 100
Num of data 10,000 5,000 10,000 1,000
Activation ReLu ReLu ReLu ReLu
Batch size 32 32 32 64
Epochs 100 200 100 50
Early stopping N/A Yes N/A N/A
We observe that the DeepSeek R1 model generates data with the greatest complexity and variety. In terms of model design, DeepSeek R1 again implements an early stopping mechanism, which was not activated in this training. Additionally, all models opt for only a few hundreds of epochs, which is usually not enough for the loss to fully converge.
Since the models are trained on different datasets, we build a universal testing set to evaluate their performances. The testing set consists of functions generated by Gaussian Random Fields (GRF) with length scales $0.05,0.1,0.2,0.5$. An example of the data is shown in Figure 7. The results for each model are shown in Table 8. Functions generated with smaller length scales are more difficult to predict for most models. The DeepSeek R1 model excelled in generalizing to unseen functions, largely due to its training data complexity. The DeepSeek V3 model failed to make meaningful predictions.
@@
Table 8: Relative $L_{2}$ errors for each DeepONet model on test datasets generated from GRF with varying length scales.
DeekSeek V3 DeepSeek R1 ChatGPT 4o ChatGPT o3-mini-high
Length scale $0.05$ 47.5% 5.7% 6.1% 6.2%
Length scale $0.1$ 51.6% 1.9% 3.4% 3.5%
Length scale $0.2$ 50.6% 0.2% 1.0% 1.0%
Length scale $0.5$ 49.4% 0.0008% 0.05% 0.06%
@@ 2.2.4 DeepONet for learning the Caputo fractional derivative operator
The input function space is specified in this test, but the LLMs are responsible for generating their own training and testing datasets. For data generation, the input function must be numerically differentiated and integrated. The methods used by each LLM for differentiation and integration are summarized in Table 9. Note that while most models implemented valid numerical schemes, ChatGPT 4o made a critical mistake in the integration step. Specifically, instead of computing the integral over the correct domain $[0,y]$, it mistakenly computed it over $[0,1]$. This led to a division by zero error as the term $(y-\tau)^{-\alpha}$ becomes singular when $y=\tau$. This bug was manually corrected and trapezoidal rule was adopted for integration.
@@
DeepSeek V3 DeepSeek R1 ChatGPT 4o ChatGPT o3-mini-high
Differentiation First-order First-order First-order Central
of $u^{\prime}$ finite difference finite difference finite difference difference
Integration Trapezoidal Left Riemann Incorrect Trapezoidal
rule (numpy) sum (from scratch) rule (numpy)
We are specifically interested in seeing if the LLMs know how to correctly encode the fractional order $\alpha$ into the DeepONet, as the fractional order $\alpha$ does not usually appear in other DeepONet applications and the LLMs’ knowledge in this particular task might be limited. In the original DeepONet paper [39], the order $\alpha$ along with the evaluation coordinates $y$ are put together into the trunk network. Notably, all four LLMs correctly placed $\alpha$ into the trunk but they repeatedly made the exact same mistake in formulating the input tensor. Specifically, they incorrectly paired each $y$ with a different $\alpha$ and structured the input as follows:
@@
However, for each fractional derivative order $\alpha$, there should be multiple evaluation points $\{y_{1},y_{2},\ldots,y_{n}\}$ at which the output function is computed. Using distinct pairs of $(y_{i},\alpha_{i})$ reduces to evaluating the fractional derivative at a single point $y_{i}$ per $\alpha_{i}$ and therefore prevents the DeepONet from learning a complete representation of the derivative across the domain on $[0,1]$. As a result, models trained solely on LLM-generated implementation, without human intervention, yield more than $150\%$ $l^{2}$ relative error. Since comparison under such an incorrect formulation does not provide meaningful insights, we manually corrected the code to ensure that the spatial input to the trunk network, $y$, is evaluated at 100 distinct points. For each function sample, the fractional order $\alpha$ is randomly chosen from the interval $[0.1,0.9]$.
@@
Additionally, DeepSeek V3 made the mistake of training the DeepONet for only a single fixed fractional order $\alpha=0.5$, even though the user request has stated clearly that $\alpha\in(0,1)$ is an arbitrary fractional order. We train and test the generated models on GRF-generated functions with length scale $0.2$ and random orders $\alpha\in(0,1)$. The results, along with model architecture and hyperparameters, are reported in Table 10.
Table 10: Network architectures, hyperparameters, and $L_{2}$ prediction errors for the four DeepONet models learning the 1D Caputo fractional derivative. The notation follows the same convention as in Table 6.
DeepSeek V3 DeepSeek R1 ChatGPT 4o ChatGPT o3-mini-high
Branch Net [100, 128, 128, 128] [100, 128, 128, 128] [100, 50, 50, 50, 100] [100, 128, 128, 100]
Trunk Net [2, 128, 128, 128] [2, 128, 128, 128] [2, 50, 50, 50, 100] [2, 128, 128, 100]
Epochs 100 50 100 100
Batch Size 32 32 32 64
Training Time (s) 11.87 6.58 13.54 10.91
Testing $L_{2}$ Error 177% 26% 19% 18%
@@
Note that the testing error for DeepSeek V3 is very high, as it was trained on a single fractional order and thus failed to generalize. Additionally, similar to our previous experiment results, all models opted for training for no more than 100 epochs, which is insufficient for proper convergence. In fact, simply increasing the number of epochs to 1000 reduces the relative error to single digits, with exactly the same model.

### v2

@@ 2.2 Scientific Machine Learning
In this section, we test the LLMs on a range of tasks in scientific machine learning, in image recognition, physics-informed machine learning, and operator learning methods.
@@ 2.2.1 MNIST digits prediction
Many problems in science and engineering require analyzing image data and classification. The MNIST [48] dataset is chosen as a test problem because it is a well-established benchmark for evaluating image recognition models. All LLMs in this experiment are asked with the following prompt:
@@
Table 5: Comparison of models generated by DeekSeek, ChatGPT, and Claude for learning the MNIST dataset. “Convolution($n$)” means a 2D convolution layer with $n$ channels, “Dense($n$)” means a fully connected layer with $n$ neurons, and “Dropout($p$)” means dropout with probability $p$.
Model Reasoning Architecture Epochs Time Testing Early
time (s) (s) Accuracy Stopping
DeekSeek V3 N/A
Convolution(32)
MaxPooling
Convolution(64)
MaxPooling
Convolution(64)
Dense(64)
Dense(10)
5 21.30 98.96% N/A
DeepSeek R1 129.0
Convolution(32)
MaxPooling
Convolution(64)
MaxPooling
Dense(128)
Dropout(0.5)
Dense(10)
20 21.23 99.11% Yes
ChatGPT 4o N/A
Convolution(32)
MaxPooling
Convolution(64)
MaxPooling
Dropout(0.5)
Dense(128)
Dense(10)
10 21.85 99.31% N/A
ChatGPT
o3-mini-high
8.0
Convolution(32)
Convolution(64)
MaxPooling
Dense(128)
Dense(10)
10 21.18 98.84% N/A
Claude 3.7 Sonnet N/A
Convolution(32)
MaxPooling
Convolution(64)
MaxPooling
Dense(64)
Dropout(0.5)
Dense(10)
10 21.30 98.98% Yes
Claude 3.7 Sonnet
extended thinking
38.0
Convolution(32)
MaxPooling
Convolution(64)
MaxPooling
Dense(64)
Dropout(0.5)
Dense(10)
10 21.33 99.14% Yes
The results show that for this relatively standard and simple problem, all models can achieve comparable training time and high testing accuracy. Both Claude models and DeepSeek R1 implemented an early stopping criterion, which may stop the training of the model early according to changes in the validation loss to prevent overfitting. Among reasoning models, ChatGPT o3-mini-high had the fastest reasoning time.
@@ 2.2.2 Physics-informed neural networks
In this section, we will test the performance of LLM in solving the PDEs with physics-informed neural networks (PINNs) [40], a framework used to solve PDEs with the power of neural networks. The Poisson equation to solve is given as
$\begin{split}-\Delta u(x,y)=1,\quad(x,y)\in[-1,1]^{2}/[0,1]^{2}\end{split}$
with zero Dirichlet boundary conditions. The question posed to all models is:
@@
All models generated code with errors. The issue was the use of the grad function in JAX to compute gradients of outputs with respect to inputs, which, when used incorrectly, can cause errors since grad is only valid for scalar output functions. After correcting this, all codes could work smoothly. The performance of the codes given by different models is summarized in Table 6.
Table 6: Network architectures, hyperparameters, and $L_{2}$ prediction errors for six PINN models solving the Poisson equation. The architecture [2,20,20,20,1] represents a fully connected neural network with three hidden layers, each containing 20 neurons, an input dimension of 2, and an output dimension of 1.
Reasoning time (s) Network Training data Epoch Training time(s) Activation function $L_{2}$ error
DeepSeek V3 N/A [2, 20, 20, 20, 1] 1000 5000 7.90 Tanh 929%
DeepSeek R1 417.0 [2, 50, 50, 50,50, 1] 512 10000 13.80 Tanh 54.7%
ChatGPT 4o N/A [2, 64, 64, 64, 64, 1] 5000 5000 15.23 Tanh 169%
ChatGPT
o3-mini-high
15.0 [2, 64, 64, 64, 1] 200 (every epoch) 5000 1740.00 Tanh 5.31%
Claude 3.7 Sonnet N/A [2, 32, 32, 32, 32,1] 2000 10000 12.10 Tanh 4.25%
Claude 3.7 Sonnet
extended thinking
38.0 [2, 32, 32, 32, 1] 5000 1000 19.00 Tanh 322%
@@
[…]
Figure 6: Predicted solutions from PINNs generated by different LLMs for solving the Poisson equation on an L-shaped domain.
Based on the results summarized in Table 6, the full reasoning model ChatGPT o3-mini-high and the hybrid reasoning model Claude 3.7 Sonnet are able to deliver correct predictions and achieve significantly lower $L_{2}$ errors compared to the other models. These results demonstrate that, when solving the Poisson equation on an L-shaped domain, reasoning models not only capture the solution’s complex behavior more accurately but also adapt more flexibly to nontrivial modifications in the problem domain.
In contrast to non-reasoning models, which typically train within the original problem domain, DeepSeek R1 also exhibits the capability to recognize and adapt to the L-shaped boundary by generating appropriate boundary points. This adaptive strategy underlines the enhanced generalization and responsiveness afforded by incorporating reasoning capabilities into the model.
An additional noteworthy observation is the difference in reasoning times. As shown in Table 6, while DeepSeek R1 requires a considerable 417 seconds of reasoning time, the other two reasoning‐based models respond much faster—ChatGPT o3-mini-high operates a reasoning time of only 15 seconds and Claude 3.7 Sonnet with extended thinking 38 seconds.
Moreover, although ChatGPT o3-mini-high requires over 1000 seconds of training time, mainly due to its need to generate new samples at every epoch, the superior accuracy it achieves justifies the additional computational effort. Overall, the findings in Table 6 provide compelling evidence that PINNs generated by reasoning LLMs offer a more flexible response to the complexities of the problem domain and significantly outperform non-reasoning models.
@@ 2.2.3 DeepONet for learning the antiderivative operator
    DeepSeek V3: Input functions are simply generated point-wise as uniformly random numbers in numpy. The functions are then numerically integrated using cumulative sums (cumsum) in numpy.
@@
    DeepSeek R1: Input functions are randomly generated from the following 4 types:
@@
      6th-degree polynomials $u(x)=c_{0}+c_{1}x+\cdots c_{6}x^{6}$ with random coefficients ranging from $[-2,2]$.
@@
      Trigonometric functions $u(x)=A\sin(\omega x)+B\cos(\omega x)$, where $A,B$ are chosen randomly from $[-2,2]$ and $\omega$ chosen randomly from $[1,5]$.
@@
      Exponential functions $u(x)=A\exp(Bx)$, where $A,B$ are chosen randomly from $[-1,1]$.
@@
      A mixture of 1, 2, and 3 given as $u(x)=c_{1}x^{2}+c_{2}\sin(\omega x)+c_{3}\exp(x)$, where $c_{i}$ are chosen randomly from $[-1,1]$ and $\omega$ from $[1,5]$. All antiderivatives are calculated analytically.
@@
    ChatGPT 4o: Input functions are generated as 5-th degree polynomial without the constant term, $u(x)=c_{1}x+\cdots c_{5}x^{5}$, where the coefficients $c_{1}$ are chosen randomly from $[-1,1]$. The antiderivatives are calculated analytically.
@@
    ChatGPT o3-mini-high: Input functions are generated as Fourier series with 3 modes $u(x)=\sum_{i=1}^{3}A_{i}\sin(2\pi ix)+B_{i}\cos(2\pi ix)$ where $A_{i},B_{i}$ are chosen randomly from $[-1,1]$. The antiderivatives are calculated numerically using cumulative trapezoidal rule with scipy.integrate.cumtrapz.
  5. 5.
    Claude 3.7 Sonnet: similar to DeepSeek R1, input functions are generated as any of the following:
    1. (a)
      Fourier series $u(x)=\sum_{i=1}^{N}A_{i}\sin(2\pi ix)+B_{i}\cos(2\pi ix)$, where $A,B$ are chosen randomly from $[-1,1]$ and $N$ chosen randomly from $\{2,3,4,5,6\}$.
    2. (b)
      Polynomial with the degree $d$ chosen randomly from $\{2,3,4,5\}$ and random coefficients from $[-1,1]$.
    3. (c)
      Trigonometric $u(x)=A\sin(Bx)+C\cos(2x)$ where $A,B,C$ are chosen randomly from $[0.5,2]$. All antiderivatives are calculated analytically.
  6. 6.
    Claude 3.7 Sonnet with extended thinking: similarly, input functions are generated as any of the following:
    1. (a)
      Polynomial with the degree $d$ chosen randomly from $\{0,1,2,3,4,5\}$ and random coefficients from $[-2,2]$.
    2. (b)
      Sine function $u(x)=A\sin(Bx)$ where $A,B$ are chosen randomly from $[0.5,3]$.
    3. (c)
      Exponential function $u(x)=A\exp(Bx)$, where $A,B$ are chosen randomly from $[0.5,2]$.
    4. (d)
      Mixed function $u(x)=A\sin(x)+Bx^{2}$, where $A,B$ are chosen randomly from $[0.5,2]$.
    5. (e)
      Gaussian $u(x)=a\exp\left(-\left(\frac{x-\mu}{\sigma}\right)^{2}\right)$ where $a\sim\text{Unif}[1,3]$, $\mu\sim\text{Unif}[-0.5,0.5]$, and $\sigma\sim\text{Unif}[0.2,0.5]$. The antiderivatives for Gaussian functions are calculated numerically using scipy.integrate.quad.
    6. (f)
      Rational function $u(x)=\frac{A}{1+Bx^{2}}$ where $a,b$ are chosen randomly from $[0.5,2]$. The antiderivatives for functions other than the Gaussian are calculated analytically.
We observe that Claude 3.7 Sonnet with extended thinking generates the most complex training dataset from a diverse mix of function types. Surprisingly, no model considers using Gaussian Random Fields (GRF) to generate training functions, which is the strategy used in the original DeepONet paper [41].
To proceed with the training, some bugs in the LLM-generated code were manually fixed. Empirically, we found through repeated experiments that Claude-generated code has more non-trivial bugs in this particular task compared to other models. A few major bugs include:
@@
    ChatGPT 4o’s code incorrectly sets the trunk net input dimension to 100 (the number of evaluation points) instead of 1 (the dimension of the dependent variable). The correct dimension was manually set and the tensor shapes were fixed accordingly.
  3. 3.
    Claude 3.7 Sonnet repeatedly outputs a bug where a KerasTensor is passed to a TensorFlow function (tf.reshape) instead of a Keras operation. This violated Keras’ Functional API rules. This bug was fixed by wrapping the function in a custom Keras layer.
  4. 4.
    Claude 3.7 Sonnet with extended thinking makes the same mistake as Claude 3.7 Sonnet.
The model architectures and training results are summarized in Table 7. Both Claude models adopt the usage of Swish activation, defined as $\text{Swish}(x)=x\cdot\frac{1}{1+e^{-x}}$, rather than the more traditional ReLu activation, justifying the choice by citing advantages like smoothness and non-monotonicity. The Claude models also used advanced training techniques such as $L^{2}$ weight regularizer, learning rate scheduling, and dropout training.
Table 7: Network architectures, hyperparameters, and $L_{2}$ prediction errors for the DeepONet models learning the antiderivative operator. The network architecture notation follows the same convention as in Table 6.
Model
Reasoning
time (s)
Bug
free ?
Branch Net Trunk Net
Num of
data
Activation
Batch
size
Epochs
Early
stopping ?
DeekSeek V3 N/A No [100, 128, 128, 128] [1, 128, 128, 128] 10,000 ReLu 32 100 No
DeepSeek R1 214.0 Yes [100,100,100,50] [1,100,100,50] 5,000 ReLu 32 200 Yes
ChatGPT 4o N/A No [100, 64, 64, 64, 64] [1, 64, 64, 64, 64] 10,000 ReLu 32 100 No
ChatGPT
o3-mini-high
11.0 Yes [100,100,100,100] [1,100,100,100] 1,000 ReLu 64 50 No
Claude 3.7 Sonnet N/A No [100, 64, 128, 64, 32] [1, 32, 64, 32, 32] 800 Swish 64 50 Yes
Claude 3.7 Sonnet
extended thinking
79.0 No [100, 256, 128, 64, 50] [1, 64, 64, 50] 2,000 Swish 128 150 Yes
Since the models are trained on different datasets, we build a universal and general testing set to evaluate their performances. The testing set consists of functions generated by Gaussian Random Fields (GRF) with length scales $0.05,0.1,0.2,0.5$. An example of the data is shown in Figure 8. The results for each model are shown in Table 8. Functions generated with smaller length scales are more difficult to predict for most models. DeepSeek R1 gives the best result, while DeepSeek V3 fails to make meaningful predictions across all length scales. We note that the poor testing performance of Claude-generated DeepONets is due to hyperparameter choices, not training data. Using the same data from Claude 3.7 Sonnet with extended thinking, we removed the regularizer, learning rate scheduler, and dropout layers, switched to ReLU activation, and increased neurons to 100 per layer to match ChatGPT o3-mini-high. With these changes, the simplified model performed comparably to both ChatGPT models. This shows that while Claude generates a good strategy, its implementation still requires human intervention and hyperparameter tuning for optimal performance.
@@ 79.0 No [100, 256, 128, 64, 50] [1, 64, 64, 50] 2,000 Swish 128 150 Yes
Table 8: Relative $L_{2}$ errors for DeepONet models on test datasets generated from GRF with different length scales. DeepSeek R1, with its diverse training data and error-free implementation, achieves the best results. Claude models used advanced dropout techniques, regularization, learning rate scheduling, and Swish activation, but without these, the models would have performed better and been comparable to ChatGPT’s results.
Model Length scale $0.5$ Length scale $0.2$ Length scale $0.1$ Length scale $0.05$
DeekSeek V3 49.4% 50.6% 51.6% 47.5%
DeepSeek R1 0.0008% 0.2% 1.9% 5.7%
ChatGPT 4o 0.05% 1.0% 3.4% 6.1%
ChatGPT o3-mini-high 0.06% 1.0% 3.5% 6.2%
Claude 3.7 Sonnet 0.35% 4.17% 12.69% 180.41%
Claude 3.7 Sonnet extended thinking 0.22% 2.05% 5.0% 7.04%
@@ 2.2.4 DeepONet for learning the Caputo fractional derivative operator
The input function space is specified in this test, but the LLMs are responsible for generating their own training and testing datasets. For data generation, the input function must be numerically differentiated and integrated. The methods used by each LLM for differentiation and integration are summarized in Table 9. Note that while most models implemented valid numerical schemes, ChatGPT 4o and Claude 3.7 Sonnet both made a similar mistake in the integration step. Specifically, instead of computing the integral over the correct domain $[0,y]$, it mistakenly computed it over $[0,1]$. This led to a division by zero error as the term $(y-\tau)^{-\alpha}$ becomes singular when $y=\tau$. This bug was manually corrected and trapezoidal rule was adopted for integration.
@@
Model Differentiation of $u^{\prime}$ Integration
DeepSeek V3 First-order finite difference Trapezoidal rule (numpy)
DeepSeek R1 First-order finite difference Left Riemann sum (from scratch)
ChatGPT 4o First-order finite difference Incorrect bounds
ChatGPT o3-mini-high Central difference Trapezoidal rule (numpy)
Claude 3.7 Sonnet Central difference Incorrect bounds
Claude 3.7 Sonnet extended thinking Central difference Trapezoidal rule (numpy)
We are specifically interested in seeing if the LLMs know how to correctly encode the fractional order $\alpha$ into the DeepONet, as the fractional order $\alpha$ does not usually appear in other DeepONet applications and the LLMs’ knowledge in this particular task might be limited. In the original DeepONet paper [41], the order $\alpha$ along with the evaluation coordinates $y$ are put together into the trunk network. Notably, all six LLMs correctly placed $\alpha$ into the trunk. However, across repeated experiments, all DeepSeek and ChatGPT models consistently made the same mistake in formulating the input tensor. Specifically, they incorrectly paired each evaluation point $y$ with a different $\alpha$ and structured the input as follows:
@@
However, for each fractional derivative order $\alpha$, there should be multiple evaluation points $\{y_{1},y_{2},\ldots,y_{n}\}$ at which the output function is computed. Using distinct pairs of $(y_{i},\alpha_{i})$ reduces to evaluating the fractional derivative at a single point $y_{i}$ per $\alpha_{i}$ and therefore prevents the DeepONet from learning a complete representation of the derivative across the domain on $[0,1]$. As a result, DeepONets trained on the original DeepSeek or ChatGPT-generated implementation, without human intervention, yield more than $150\%$ $l^{2}$ relative error when tested. Since comparison under such an incorrect formulation does not provide meaningful insights, we manually corrected the code by DeepSeek and ChatGPT to ensure that the spatial input to the trunk network, $y$, is evaluated at 100 distinct points per the fractional order $\alpha$, which is randomly chosen from the interval $[0.1,0.9]$.
Both Claude models correctly formulated the trunk net input in all experiments. However, they consistently tiled the branch net and trunk net input tensors to match their first dimension before passing them into the network. This ensured that their dot product could be directly computed after they passed through the two branches, but this tiling was unnecessary and expensive, as the dot product could be implemented easily with einsum or matrix multiplication in TensorFlow. The excessive tiling led to very large tensor sizes and slowed down training.
Additionally, DeepSeek V3 mistakenly trained the DeepONet using only a fixed fractional order $\alpha=0.5$, despite the user request specifying $\alpha\in(0,1)$ as arbitrary. Similarly, Claude 3.7 Sonnet with extended thinking used five fixed fractional orders, $\alpha=0.17,0.25,0.34,0.65,0.73$, across all 900 training functions, resulting in a total training sample size of 4500. Fixing the fractional order in training data reduced the DeepONet’s ability to generalize to unseen fractional orders during testing. The correct approach is to sample $\alpha$ randomly for each training sample. However, these mistakes were not manually corrected as they are part of the LLMs’ decision making.
@@
Finally, we train and test all DeepONets on GRF-generated functions with length scale $0.2$ and random orders $\alpha\in[0.1,0.9]$. The results, along with model architecture and hyperparameters, are reported in Table 10.
Table 10: Network architectures, hyperparameters, and $L_{2}$ prediction errors for the DeepONet models learning the 1D Caputo fractional derivative. The notation follows Table 6. The testing $L_{2}$ errors reported in this table reflect corrections made after manually fixing bugs in the original code. Claude 3.7 Sonnet achieves the best result but is slower due to inefficient tensor manipulation in the code. The performances of DeepSeek V3 and Claude 3.7 Sonnet with extended thinking are hindered by using a fixing fractional order $\alpha$ across different training samples.
Model
Reasoning
Time (s)
Bug
free ?
Branch Net Trunk Net Epochs
Num of
train data
Activation Training Time (s) Testing $L_{2}$ Error
DeepSeek V3 N/A No [100, 128, 128, 128] [2, 128, 128, 128] 100 1000 ReLu 11.87 177.19%
DeepSeek R1 426.0 No [100, 128, 128, 128] [2, 128, 128, 128] 50 1000 ReLu 6.58 26.82%
ChatGPT 4o N/A No [100, 50, 50, 50, 100] [2, 50, 50, 50, 100] 100 1000 ReLu 13.54 19.49%
ChatGPT
o3-mini-high
23.0 No [100, 128, 128, 100] [2, 128, 128, 100] 100 1000 ReLu 10.91 18.41%
Claude 3.7 Sonnet N/A No [100, 128, 128, 40] [2, 128, 128, 40] 200 900 ReLu 41.74 5.59%
Claude 3.7 Sonnet
extended thinking
109.0 No [100, 128, 128, 128, 50] [2, 128, 128, 128, 50] 50 4500 ReLu 1943.14 16.24%
@@
The testing $L_{2}$ errors reported in this table reflect corrections made after manually fixing bugs in the original code. Note that the testing error for DeepSeek V3 is very high, as it is trained on a single fractional order and thus fails to generalize. Claude 3.7 Sonnet with extended thinking also suffers from the same mistake, but its larger sample size and choice of five different $\alpha$’s contributes to better performance. Overall, the hybrid reasoning model Claude 3.7 Sonnet achieves the best result but is slower due to inefficient tensor manipulation in the code. Additionally, similar to our previous experiment results, most models are only trained for no more than 100 epochs, which is insufficient for proper convergence. In fact, during our experiments, simply increasing the number of epochs to 1000 reduces the relative error to single digits, with exactly the same model.